OpenStudy (anonymous):

Integral of 13y/(y^4+1). Can someone explain the process of solving this?

ganeshie8 (ganeshie8):

sub \(u =y^2\)

OpenStudy (raden):

use trigo sub : y^2 = tan(theta)

OpenStudy (raden):

y^2 = tanθ 2y dy = sec^2 (θ) dθ so, int 13y/(y^4 +1) dy = 13/2 int (sec^2 θ)/(tan^2 θ + 1) dθ = 13/2 int (sec^2 θ)/(sec^2 θ) dθ = 13/2 int dθ = ....

OpenStudy (anonymous):

ahhhh okay, thanks! :)

OpenStudy (raden):

welcome :)